OpenStudy (anonymous):

ATCHING EXERCISE Match the terms with the definitions. What is the slope of a line parallel to 4x + 2y = 5? What is the slope of a line perpendicular to x - 3y = 9? What is the slope of a line parallel to 10x - 5y = 8? What is the slope of a line perpendicular to x - 5y = -10? What is the slope of a line parallel to 4x + y = -1?

OpenStudy (anonymous):

A) m = 2 B) m = -5 C) m = -2 D) m = -3 E) m = -4

OpenStudy (cwrw238):

A m = -2

OpenStudy (anonymous):

huh its matching..

OpenStudy (anonymous):

What is the slope of a line parallel to 4x + 2y = 5? What is the slope of a line perpendicular to x - 3y = 9? What is the slope of a line parallel to 10x - 5y = 8? What is the slope of a line perpendicular to x - 5y = -10? What is the slope of a line parallel to 4x + y = -1? you use those and match with the number 1,2,3,4,5

OpenStudy (cwrw238):

1 = C then? right?

OpenStudy (anonymous):

yes sorry i got confused

OpenStudy (anonymous):

all those equations are standard form: \(\large Ax+By=C \) the slope is obtained by: \(\large m=-\frac{A}{B} \) calculate the slope using that and you can do your matching easily...